### v1

# An Experimental Evaluation of Imputation Models for Spatial-Temporal Traffic Data

Abstract
Traffic data imputation is a critical preprocessing step in intelligent transportation systems, enabling advanced transportation services. Despite significant advancements in this field, selecting the most suitable model for practical applications remains challenging due to three key issues: 1) incomprehensive consideration of missing patterns that describe how data loss along spatial and temporal dimensions, 2) the lack of test on standardized datasets, and 3) insufficient evaluations. To this end, we first propose practice-oriented taxonomies for missing patterns and imputation models, systematically identifying all possible forms of real-world traffic data loss and analyzing the characteristics of existing models. Furthermore, we introduce a unified benchmarking pipeline to comprehensively evaluate 10 representative models across various missing patterns and rates. This work aims to provide a holistic understanding of traffic data imputation research and serve as a practical guideline.
Index Terms:
traffic data imputation, spatial-temporal graph series, experimental evaluation.
I Introduction
Traffic data is often represented with a graph of spatial-temporal series, where each node in the graph represents a traffic sensor continuously collecting observation series and edges in the graph describe the relationships between sensors. Traffic data holds a prominent role in Intelligent Transportation Systems (ITS), supporting key services in ITS, such as real-time traffic display, traffic data prediction, and traffic signal control. Despite the increasing deployment of sensors to collect real-time traffic information, missing data remains a persistent issue due to various factors like equipment malfunctions, power outages, and network failures. These missing data hinder ITS to accurately perceive traffic changes over time, resulting in the performance decline of transportation services. Therefore, traffic data imputation is a crucial preprocessing step in ITS, ensuring the quality of traffic data.
Numerous models have been proposed to solve this problem, which primarily focus on addressing two challenges: 1) Effectively capturing the dynamics of traffic data with missing values across both spatial and temporal dimensions. 2) Training models with the absence of ground truth for the missing data. Beyond specialized traffic imputation models, general multivariate time series imputation models can also be adapted for this problem. Despite significant progress in this field, choosing an optimal model for specific real-world scenarios in ITS remains challenging due to three factors:
    Missing patterns are not comprehensive. In real-world scenarios, missing patterns in traffic data, which reflect how data loss along the spatial and temporal dimensions, exhibit significant diversity. For example, network failures may cause random traffic data loss, sensor malfunctions can result in continuous traffic data loss in the temporal dimension, and power outages may lead to sustained traffic data loss across both spatial and temporal dimensions. However, most existing studies consider only a limited subset of these patterns, leaving the performance of their models on real-world traffic data with other missing patterns largely unexplored.
    Datasets are not standardized. Existing studies often test their imputation models on different datasets, with model designs tailored to the specific characteristics of each datasets. There is a significant lack of empirical studies comparing different models on a standardized set of datasets, especially for spatial-temporal traffic data. Additionally, while both traffic flow and traffic speed are critical for ITS, few studies have conducted comprehensive experiments on both simultaneously.
    Evaluations are not sufficient. Firstly, the methods evaluated within a single paper are often limited in scope, e.g., newly proposed approaches from the same period are rarely compared empirically. Secondly, although traffic prediction models can also be adapted for traffic data imputation, systematic comparisons between imputation and prediction models under consistent settings are scarce. Thirdly, most evaluations only report overall average errors, neglecting performance analysis during difficult periods with significant traffic fluctuations, which obscures differences in model performance under such scenarios. Thirdly, model efficiency analysis, a key factor for real-world deployment, is frequently neglected.
To meet the need, we design an unified benchmarking pipeline to evaluate ten representative or recently proposed imputation models for spatial-temporal traffic data. Our evaluation spans both traffic flow and speed datasets, covering all possible missing patterns across spatial and temporal dimensions, accompanied by detailed performance analysis. Meanwhile, to provide readers with an overall picture of the latest advancements on traffic data imputation and to enhance their insights into model characteristics, we offer a holistic overview of imputation models for traffic data and categorize various methods from a practice-oriented perspective.
While there are several recent surveys on data imputation models, they pay little attention to their applications in spatial-temporal traffic scenarios. For instance, Miao et al. [1] provides a general discussion on imputation across various data types (including numerical, categorical, and mixed data). Fang et al. [2] focuses on introducing deep learning-based imputation models for time series without empirical evaluations. Wang et al. and Du et al. review deep learning methods for multivariate time series imputation and introduce benchmark platforms [3, 4], yet they exclude imputation approaches beyond deep neural networks, such as tensor completion methods. Compared to them, this paper distinguishes itself in two key aspects. Firstly, we emphasize representative and emerging imputation models specifically applicable for spatial-temporal traffic data. Secondly, beyond simply analyzing overall experimental results, we provide visualizations and delve into a detailed performance analysis during critical difficult periods, addressing core concerns within the ITS domain. The contributions of this work are summarized as follows:
    Practice-oriented taxonomy. Focusing on the task of spatial-temporal traffic data imputation, we present taxonomies for both missing patterns and imputation models. Missing patterns are categorized into four types, each representing a distinct form of real-world data loss. Imputation models are classified based on two key aspects of practical code implementation: spatial-temporal modeling techniques and loss function design.
    Unified benchmarking pipeline. We establish a pipeline to test imputation models for spatial-temporal traffic data. The pipeline unifies the test process including missing scenarios construction, data pre-procesing, model construction and performance evaluation. Implementation of the pipeline is publicly available at https://github.com/wtl52656/imputation_benchmark.
    Comprehensive evaluation. We conduct extensive experiments on 10 representative models across three traffic datasets, evaluating their performance under 20 scenarios (comprising combinations of 4 missing patterns and 5 missing rates). A detailed analysis of both effectiveness and efficiency is provided to serve as a practical guideline for model selection and application.
The rest of this paper is organized as follows: Section II reviews related work. Section III gives an introduction of the spatial-temporal traffic data imputation problem. Section IV presents a overview on imputation models and proposes the taxonomy. Section V introduces the evaluated models. Section VI presents the unified benchmarking pipeline and experimental results. Section VII concludes the paper.
II Related Work
In this section, we review four categories of works relevant to spatial-temporal traffic data imputation. Specifically, multi-variable time series imputation is a broad topic, with spatial-temporal graph data imputation as its subbranch. Spatial-temporal graph imputation models are the most suitable kind of approaches to achieve traffic data imputation. Besides, spatial-temporal kriging is a special case of spatial-temporal graph imputation when there are continuous missing data in the temporal dimension. Time series probabilistic imputation further incorporates uncertainty for the imputation values.
II-A Muti-variable Time Series Imputation
Multivariate time series imputation (MTSI) has been a widely researched topic over the years, focusing on the exploitation of both temporal dependencies and feature relationships to fill in missing data. Early works utilize statistical techniques, such as the last observed value or employing k-nearest neighbors algorithms [5, 6]. Tensor completion (TC) is another branch of MTSI, which harnesses the assumption of low-rank matrices to represent global correlations to impute missing data, with representative works including [7, 8, 9, 10, 11, 12, 13, 14].
Within the domain of deep learning, many MTSI works have been proposed [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], predominantly relying on recurrent neural networks (RNNs) and Transformers to capture temporal dependence. To make the imputed data more realistic, some methods utilize adversarial training strategy [34, 35, 36, 37, 38, 39, 40, 41]. To capture both local and global temporal correlations, LGnet [42] combines LSTMs for local dynamics with memory networks for global dynamics. NAOMI [43] models long-range dependencies using a multiresolution structure that recursively imputes missing values from coarse to fine resolutions via a divide-and-conquer strategy. mTAN [44] enhances the capture of local structures by applying the attention mechanism to latent RNN states. PrimeNet [45] introduces an attention-based pretraining method that integrates time-sensitive contrastive learning for data imputation.
II-B Spatial-temporal Graph Imputation
In contrast to MTSI, the spatial-temporal graph imputation (STGI) approach [46, 47, 48, 49, 50] incorporates an extra graph structure to explicitly represent the spatial correlations inherent in time series data. Typically, STGI utilizes recurrent neural networks (RNNs) and attention mechanisms to capture temporal dependencies, combined with graph neural networks (GNNs) to capture spatial dependencies. For example, GRIN [51] utilizes a specialized message-passing GNN to effectively capture spatial dependencies and impute time series in both forward and backward directions. SPIN [52] incorporates sparse spatial-temporal attention mechanisms, which only propagate observed data points, thus efficiently diminishing noise associated with missing values. MDGCN [53] strengthens the modeling of spatial correlations by learning the graph structure and integrates an external memory network to retain global spatiotemporal knowledge. GCASTN [54] introduces a generative-contrastive self-supervised learning framework, enhancing model robustness by comparing two different augmented perspectives. ImputeFormer [55] presents a Transformer-based architecture that leverages low-rankness principles to bridge the gap between deep learning and tensor completion methods for data imputation.
II-C Spatial-temporal Kriging
Spatial-temporal Kriging (STKriging) is designed to impute data at locations where sensors are not deployed [56, 57, 58, 59, 60], meaning the complete historical time series for these locations is unavailable. To address this, the common approach is leveraging surrounding sensors for imputation. KCN [61] is the first deep learning-based approach for STKriging, utilizing graph neural networks (GNNs) and k-nearest neighbor models to impute unobserved locations by modeling spatial correlations. Building on this foundation, IGNNK [62] introduces a message-passing mechanism within GNNs. It is the first work to impute missing data in an inductive manner. Subsequently, many methods follow this line. For example, SAGCN[63] integrates temporal dependencies using temporal convolutional networks (TCN) and spatial dependencies through a multi-aggregator mechanism. INCREASE[64] incorporates heterogeneous spatial relations from three perspectives and leverages gated recurrent units (GRU) to model temporal correlations. DualSTN[65] introduces a skip graph GRU to capture long-term spatial-temporal correlations. IAGCN[66] learns an adaptive graph structure to reconstruct spatial relationships, and STGNP [67] incorporates uncertainty estimates for imputation. ESC-GAN [68] partitions the interpolated area into grids, modeling local correlations with convolutional neural networks (CNNs) and employing a generative adversarial network (GAN) to train the model effectively.
II-D Time Series Probabilistic Imputation
The imputation above methods are focused on deterministic imputation, where only a plausible value is filled in for the missing data. In contrast, Time Series Probabilistic Imputation (TSPI) seeks to provide a range for the missing values, thereby quantifying the confidence level in the imputation [35]. A typical approach in this domain involves using generative models to capture the data distribution. For instance, GP-VAE [69] utilizes a Gaussian process as a prior to represent temporal dependencies.
Recently, diffusion models have emerged as a promising technology for TSPI. For example, CSDI [70] harnesses the stochastic characteristics of diffusion models to estimate uncertainty. It processes observed data and imputes missing values through a multi-step denoising procedure. PriSTI [71] derives prior information from the available data and employs spatiotemporal attention for imputation. FastSTI [72] incorporates a high-order pseudo-numerical solver to enhance inference speed. MTSCI [73] integrates a contrastive complementary mask strategy and a mixup technique into the diffusion model, leveraging conditional information from neighboring windows to ensure consistent imputation.
III Preliminaries
In this section, we first formulate the spatial-temporal traffic data imputation problem, then present a taxonomy on missing patterns in traffic data.
III-A Problem Statement
Definition 1.
[…]
Definition 2.
Traffic Data. The traffic data observed on node $v\in V$ at time slice $t$ is denoted as $x_{t,v}\in\mathbb{R}$. The whole traffic data observed on $G$ at time slice $t$ are defined as $\mathbf{x}_{t}=(x_{1,t},x_{2,t},\cdots,x_{N,t})\in\mathbb{R}^{N}$. And we use $\mathbf{X}=(\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{T})\in\mathbb{R}^{N\times T}$ to denote all observations over $T$ time slices.
Definition 3.
Masking Matrix. To indicate the missing position in the observed traffic data, it is necessary to introduce an observation masking matrix $\mathbf{M}\in\mathbb{R}^{N\times T}$, where $m_{v,t}=0$ when $x_{v,t}$ is missing, and $m_{v,t}=1$ when $x_{v,t}$ is observed.
Definition 4.
Time Lag Matrix. To record the time lag between current traffic data and the last observed traffic data, a time lag matrix $\boldsymbol{\Delta}\in\mathbb{R}^{N\times T}$ is usually introduced. Each element $\delta_{v,t}\in\boldsymbol{\Delta}$ is defined as follows,
$\delta_{v,t}=\begin{cases}0&\text{ if }t=1\\ 1&\text{ if }t>1\text{ and }m_{v,t-1}=1\\ \delta_{v,t-1}+1&\text{ if }t>1\text{ and }m_{v,t-1}=0\end{cases}$ (1)
Definition 5.
Spatial-Temporal Traffic Data Imputation. Given the incomplete observed traffic data $\mathbf{X}$ and the corresponding traffic network $G$ and masking matrix $\mathbf{M}$, the goal is to estimate complete traffic data $\hat{\mathbf{X}}$.
III-B Missing Patterns
Missing patterns denote how the traffic data miss. Different missing patterns correspond to various types of real-world scenarios. This paper classifies the missing patterns in traffic data into four categories depending on the dimensions where missing happens (e.g., the spatial or temporal dimension) and whether the missing values appear continuously or randomly. Specifically, Fig. 1 illustrates these four missing patterns, including SRTR, SRTC, SCTR, and SCTC, where S/T indicates the Spatial/Temporal dimension, and R/C indicates missing positions are random/continuous.
Figure 1: Illustrations of four missing patterns in traffic data, where white grid indicates missing data and gray grid indicates observed data. Specifically, SRTR may be caused by random signal or network interruption, SCTR may be caused by equipment failure over a group of devices due to some factors like network interruption in an area, SRTC may be caused by some sensors experiencing equipment failure, or network outages over a period of time, and SCTC may be caused by some reasons like power failure on a group of devices over a period of time in an area.
IV Models Summary
In this section, we start by providing a concise overview on models for time series imputation, within which traffic data imputation is a special case. Next, we introduce a practice-oriented taxonomy for traffic data imputation models, emphasizing their commonalities.
IV-A Overview on Imputation Models for Time series
To provide a overall picture on research popularity for time series imputation models, we present statistics on the annual number of related published papers, as shown in Figure 2. We also summarize the imputation model on key techniques, training methods, tasks, and dataset as shown in Table I. Specifically, we search for imputation-related papers in the IEEE Xplore, ACM digital libraries, and Web of Science. Our statistics consider the papers published from 2018 to 2024 in prestigious venues, e.g., renowned conferences and journals, including NeurIPS, ICLR, AAAI, ICML, IJCAI, ICDE, VLDB, KDD, TKDE, TITS. To ensure the papers focus on time series imputation, we use keywords including“missing data,” “imputation data,” and “interpolation” during our search. Finally, based on these papers, we conduct a manual filtering process to select the papers on time series imputation.
Figure 2: The annual number of papers published on time series imputation from 2018 to 2024.
TABLE I: Summary on the Imputation Models.
Reference Model Tasks Key Techniques Training Method Dataset Cites
$\bigstar$ Cao et al. - NeurIPS’18 [74] BRITS MTS imputation RNN Self-supervised Air Quality/ Health-care/ Human Activity 426
Shao et al. - TITS’18 [11] CTD STG imputation MF self-supervised Traffic 18
Kim et al. - IJCAI’18 [15] TBF MTS imputation RNN self-supervised Health-care 65
Luo et al. - NeurIPS’18 [18] GRUI-GAN MTS imputation RNN GAN Air Quality/ Health-care 402
Yoon et al. - ICML’18 [35] GAIN Probabilistic imputation GAN GAN UCI 859
Che et al. - ICML’18 [19] MR-HDMM MTS imputation RNN self-supervised Health-care/Air Quality 43
Yoon et al. - ICLR’18 [20] M-RNN MTS imputation RNN self-supervised Health-care 30
Wang et al. - TITS’19 [56] OCC ST Kriging RNN self-supervised Traffic 18
Li et al. - TITS’19 [48] MVLM STG imputation RNN self-supervised Traffic 151
Wang et al. - TITS’19 [12] TAS-LR STG imputation MF unsupervised Traffic 73
$\bigstar$ Luo et al. - IJCAI’19 [34] E2GAN Probabilistic imputation RNN GAN Air Quality/Health-care 166
Liu et al. - NeurIPS’19 [43] NAOMI Probabilistic imputation RNN GAN Traffic 102
Hwang et al. - ICML’19 [37] Hexagan Probabilistic imputation GAN GAN/semi-supervised Images/Synthetic 44
Li et al. - ICLR’19 [36] Misgan Probabilistic imputation GAN GAN Images 192
Ma et al. - TKDE’19 [21] OSICM MTS imputation kNN self-superviesd Air Quality/Health-care/Wiki4HE 13
Zhang et al. - ICDE’19 [22] IIM MTS imputation KNN self-supervised UCI/KEEL/Siemens 37
Wu et al. - ICDE’19 [23] HKMF-T MTS imputation MF self-supervised Traffic 8
Park et al. - KDD’19 [38] Imp-GAIN MTS imputation RNN+Attention GAN Health-care 12
Jiang et al. - ICML’20 [24] BiLSTM-A MTS imputation RNN self-supervised Air Quality 1
Khayati et al. - VLDB’20 [7] ORBITS MTS imputation Centroid Decomposition self-supervised Soccer/MotionSense/BAFU/Gas 5
Ma et al. - KDD’20 [25] MIDIA MTS imputation AutoEncoder self-supervised Air Quality/Adult/Car 17
$\bigstar$ Shukla et al. - ICLR’21 [44] mTAN MTS imputation Attention supervised Clinical /Human Activity 86
Gong et al. - IJCAI’21 [13] SMV-NMF STG imputation k-means + MF unsupervised Urban Statistical 20
Xiao et al. -TITS’21 [39] MTCIU GAN MTS imputation GAN GAN Traffic 3
$\bigstar$ Chen et al. - TITS’21 [14] LATC MTS imputation TC unsupervsed Traffic 47
Zhang et al. - TITS’21 [40] SA-GAIN MTS imputation Attention GAN Traffic 27
Deng et al. - TITS’21 [49] GTC STG imputation TC unsupervsed Traffic 16
Xu et al. - TITS’21 [26] GA-GAN Probabilistic imputation GraphSAGE GAN Traffic 5
Miao et al. - AAAI’21 [27] SSGAN MTS imputation RNN semi-supervised Human activity/Clinic/Meteorologic 54
Tashiro et al. - NIPS’21 [70] CSDI Probabilistic imputation Attention + Diffusion Model self-supervised Air Quality / Health 118
Dai et al. - ICML’21 [41] MI-GAN Probabilistic imputation GAN GAN ADNI/Synthetic 11
Qin et al. - KDD’21 [46] ST-SCL STG imputation VAE semi-supervised Traffic 16
Liang W et al. - TITS’22 [57] STAR ST Kriging GNN+Attention supervised Traffic 50
Lei M et al. - TITS’22 [58] BKMF ST Kriging GP unsupervised Health-care 11
Wang A et al. - TITS’22 [47] GSTAE STG Imputation GCN+GRU supervised Traffic 0
Li J et al. - TKDE’22 [17] MT-CSR MTS Imputation CNN supervised Traffic 2
Luo X et al. - TKDE’22 [28] NeuLFT MTS Imputation TC unsupervised Dynamic interaction networks 63
Zhao Y et al. - AAAI’22 [29] GC MTS Imputation EM unsupervised Synthetic 6
Fan J. - AAAI’22 [10] D-NLMC MTS Imputation TC unsupervised Synthetic Air Quality / Temperature / Chlorine Level Dataset 1
Marisca I et al. - NIPS’22 [52] SPIN MTS Imputation Attention supervised Traffic 9
Ipsen N B et al. - ICLR’22 [75] supMIWAE Probabilistic Imputation Autoencoder supervised IMAGE / CLASSIFICATION 29
Liu Y et al. - ICLR’22 [30] MBMF Probabilitisc Imputation MissForest supervised Synthetic / Medical 3
Lao D et al. - KDD’22 [59] WSGNN ST Kriging GCN weakly-supervised Cora / Citeseer / Pubmed 8
Xing J et al. - TITS’23 [8] DFCP MTS imputation TC unsupervised CL / LPR 0
Xiuqin Xu et al. - TITS’23 [9] HRST-LR MTS imputation TC unsupervised Traffic 0
Varga B et al. - TITS’23 [60] STK ST Kriging Kriging unsupervised Traffic 1
Blázquez-García A et al. - TKDE’23 [31] MGP MTS imputation GP unsupervised Synthetic / CLASSIFICATION 0
Chowdhury R R et al. - AAAI’23 [45] PrimeNet MTS imputation Attention self-supervised PhysioNet / MIMIC-III 0
Zhang X et al. - ICML’23 [16] UTDE-mTAND MTS imputation Attention supervised MIMIC-III 1
Zhao H et al. - ICML’23 [32] TDM MTS imputation NN unsupervised UCI 10
Li X et al. - ICDE’23 [33] BiSIM MTS imputation RNN + Attention unsupervised Indoor Positioning 0
$\bigstar$ Li M et al. - ICDE’23 [71] PriSTI MTS imputation GNN + Attention + Diffusion Model self-supervised Air Quality / Traffic 4
Liu H et al. - VLDB’23 [50] HMTRL STG imputation GCN + GRU + Attention self-supervised Traffic 4
Chen P et al. - TITS’24 [76] LRTC-SCAD MTS imputation TC unsupervised Traffic 0
Zeng Z et al. - TITS’24 [77] TNN-HTV MTS imputation TC unsupervised Traffic 2
Chen P et al. - TITS’24 [78] T-DAE STG imputation Transfer Learning GAN Traffic 2
Yang H et al. - TITS’24 [79] LFA-TRCE MTS imputation TC unsupervised Traffic 0
Shu H et al. - TITS’24 [80] LRTC-3DST STG imputation TC unsupervised Traffic 1
Wei X et al. - TITS’24 [81] SAGCIN STG imputation Attention + GCN self-supervised Traffic 0
Li B et al. - TITS’24 [82] CLRTR MTS imputation TC unsupervised Traffic 2
Cheng S et al. - TITS’24 [72] FastSTI STG imputation Attention + GCN + Diffusion Model self-supervised Traffic 0
Liu D et al. - TKDE’24 [83] SGMCAI-DiT MTS imputation Attention + Diffusion Model self-supervised Industry 5
Chen X et al. - TKDE’24 [84] LCR-2D MTS imputation TC unsupervised Traffic 17
Park Byoungwoo et al. - NeurIPS’24 [85] PBDF MTS imputation Attention self-supervised Weather / Clinical 0
Obata Kohei et al. - KDD’24 [86] MissNet MTS imputation State-Space Model self-supervised Synthetic / Motion / Temperature 0
Nie T et al. - KDD’24 [55] ImputeFormer STG imputation Attention self-supervised Traffic 13
Zhang Z et al. - KDD’24 [87] LTVTI MTS imputation Attention self-supervised AIS 0
IV-B Model Taxonomy for Traffic Data Imputation
To deepen our understanding of the similarities and differences among various imputation models for traffic data, we propose a practice-oriented taxonomy based on two criteria: spatial-temporal modeling techniques employed to capture the correlations in data, and loss design utilized for model training. Both criteria are closely tied to practical code implementation. Notably, an imputation model for traffic data may integrate a combination of these techniques and design its loss according to its specific model type and training strategy.
IV-B1 Spatial-Temporal Modeling Techniques
The key point in traffic data imputation is to leverage the spatial-temporal correlation of data to infer the missing values. So, in this subsection, we briefly introduce the widely employed techniques for spatial and temporal correlation learning in imputation task, including recurrent neural networks (RNN), attention mechanism, graph neural networks (GNN) [88], and tensor completion (TC) [89].
    RNN is a sequential model commonly used to process time series data. It operates by applying the same transformation at each time step, taking the input of the current time step and the hidden state from the previous step to produce the hidden state for the current step. This sequential approach captures the temporal evolution of traffic data. Advanced variants, such as Long Short-Term Memory (LSTM) [90] and Gated Recurrent Unit (GRU) [91], are designed to capture long-term temporal dependencies.
    Attention offers a global receptive field, enabling the modeling of correlations between elements in a sequence regardless of their distance. This flexibility allows it to effectively capture the complex dynamics of traffic data and addresses the long-term dependency issue that RNNs struggle with. However, its quadratic space complexity demands significant computational resources.
    GNN is widely used to capture the spatial correlations in graph-structured traffic data. The core design of GNNs is the pairwise message-passing mechanism, which enables nodes in the graph to iteratively update their representations by exchanging information with neighboring nodes. Representative GNN includes Graph Convolutional Network (GCN) [92], Graph attention network (GAT) [93].
    TC is a technique used to recover missing entries in a tensor. Matrix Completion (MC) is a special case of TC when the tensor has only two dimensions. TC based imputation models typically assume that the underlying tensor has a low-rank structure, optimizing it through rank minimization. This approach enables the models to capture global patterns, such as daily and weekly periodicity, in traffic data.
IV-B2 Loss Design
The design of the loss function for imputation model training is guided by two factors: model types and training strategies.
Regarding model types, there are two primary categories: predictive models and generative models. Predictive imputation models focus on learning the direct relationship between conditioned observed traffic data and the imputation target. The loss functions used for these models, such as Mean Absolute Error (MAE) and Mean Squared Error (MSE), measure the distance between the predicted output and the ground truth. On the other hand, generative imputation models aim to learn the joint probability distribution of the conditioned observations and the imputation target or to generate new data samples that align with given conditions. The loss functions for generative imputation models vary with the specific type of generative model used. Here, we introduce the most common generative models for time series imputation, including Generative Adversarial Networks (GANs) [94], Variational Autoencoders (VAEs) [95], and diffusion models [96, 97, 98].
    VAE is a generative model combining deep learning and Bayesian inference. It consists of an encoder and a decoder, like a traditional autoencoder. In particular, it introduces a probabilistic latent space, encoding inputs as distributions (typically Gaussian) rather than fixed points. Such probabilistic nature of VAEs allows them to model uncertainty. The VAE loss function, derived from the variational lower bound, consists of two parts, the reconstruction loss that measures how well the decoder reconstructs the input data from the latent space, and the KL divergence loss that ensures that the learned latent distribution stays close to a prior distribution. Formally,
    $\begin{split}\mathcal{L}=\mathbb{E}_{q(z|x)}[\mathrm{log}\,p(x|z)]-D_{KL}(q(z|x)||p(z)),\end{split}$ (2)
    where $z$ is the latent variable and $p(z)$ is its prior distribution. $q(z|x)$ is the learned latent distribution. $p(x|z)$ represents the likelihood of the observed data $x$ given a latent variable $z$.
    GAN is composed of two main components: a generator $G$ and a discriminator $D$. The generator aims to transform random noise $z$ into fake data. Within the domain of data imputation, incorporating observed data into the generator’s input is crucial to prevent the generation of biased imputation target. The discriminator receives both the fake data produced by the generator and truth data, then aims to assess the likelihood of each being genuine. Through adversarial training, the generator enhances its capability to produce data that closely mimics the real thing, while the discriminator strengthens its proficiency in distinguishing between authentic and synthetic data. The two components play a min-max game and the loss function is:
    $\small\mathcal{L}=\min_{G}\max_{D}E_{x\sim q(x)}[\mathrm{log}\ D(x)]+E_{z\sim p(z)}[\mathrm{log}(1-D(G(z)))]$ (3)
    Diffusion model is a type of deep generative model. It begins by adding random noise to data samples over several time steps, progressively transforming them into noise, and then learns to reverse this process to recover the original data. During training, the model learns to denoise noisy data step by step, making it capable of generating new, realistic, structured samples. The training objective for diffusion models typically aims to minimize the difference between the model’s predicted denoised sample and the actual clean sample at each diffusion step. In practice, the training loss is often formulated as a mean squared error (MSE) that measures the distance between the predicted noise and the actual noise added at each step in the forward diffusion process:
    $\mathcal{L}=||\epsilon_{t}-\epsilon_{\theta}(x_{t},t)||^{2},$ (4)
    where $\epsilon_{\theta}(\cdot)$ represents the denoising network, which predicts the amount of noise that needs to be removed, and $\epsilon_{t}\sim\mathcal{N}(0,1)$ is the actual noise added.
Regarding training strategy, a critical consideration in real-world scenarios is the absence of ground truth for missing data, as past missing data cannot be retrieved. Consequently, imputation models cannot be trained against the actual ground truth. To address this, self-supervised learning (SSL) is employed. SSL methods for imputation can be further categorized by the way how observed data is used to guide model training: reconstruction-based SSL which reconstructs observed data to guide model training, and masked SSL which trains models by first masking portions of the observed data and then recovering them. Next, we delve these two training strategies.
    Reconstruction-based SSL strategy guides the training of imputation model by reconstructing observed data. The corresponding loss function is usually defined as follow:
    $\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}|\mathbf{X}_{i}\odot\mathbf{M}-\hat{\mathbf{X}}_{i}\odot\mathbf{M}|,$ (5)
    where $i$ is the sample index. Broadly, two types of models usually adopt this training strategy, autoencoder-based imputation models and TC-based imputation models. Specifically, autoencoder-based imputation models reconstruct the observed data by passing it through a bottleneck layer, learning compressed representations that capture meaningful patterns in the data. On the other hand, TC-based imputation models reconstruct the observed data under the assumption of a low-rank structure, effectively capturing global patterns and dependencies.
    Masked SSL strategy for training imputation models is inspired by masked language modeling [99]. Given a sample $\mathbf{X}$, the observed values in $\mathbf{X}$ are divided into two parts, then one part of them is set as the imputation targets $\mathbf{X}^{ta}$ and the other one part is set as the conditional observations $\mathbf{X}^{co}$. When training the imputation model, for each sample, the conditional observations $\mathbf{X}^{ta}$ is the input, and the imputation target $\mathbf{X}^{ta}$ is the deserved output. Finally, the imputation is trained by comparing the reconstruction error with regard to the imputation target $\mathbf{X}^{ta}$. In practice, MAE is commonly chosen as the loss function. Formally,
    $\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}|\mathbf{X}_{i}^{ta}-\hat{\mathbf{X}}_{i}^{ta}|,$ (6)
    where $i$ is the sample index. Masked self-supervised learning is able to simulate various missing scenarios during training by changing the choice of imputation target. Thus, the choice of imputation target is a key influencing the effectiveness of imputation in real-world application.
V Models for evaluation
In this section, we briefly introduce the models evaluated in this paper. Specifically, we establish three criteria for model selection. 1) Papers that focus on traffic data imputation. 2) Papers addressing multivariate time series imputation with high citation counts, as traffic data can be treated as a special type of multivariate time series. 3) Papers on traffic data prediction, as both prediction and imputation tasks rely on observed data to infer unobserved data. Prediction models can be adapted for imputation by modifying their training and data processing methods. Based on these criteria, we select the following 10 models for evaluation.
V-A BRITS
BRITS [74] is a bidirectional RNN-based model for irregular multivariate time series imputation. It is one of the earliest representatives that introduce data-driven deep-learning technologies to release existing imputation models’ strong specific assumptions on the underlying data-generating process. BRITS uses a “complement” input estimated by historical observations and other features at the same time $\mathbf{c}_{t}$ when $\mathbf{x}_{t}$ is missing. Then, it introduces a temporal decay factor $\gamma_{t}=exp\left\{-\max(0,\textbf{W}_{\gamma}\delta_{t}+\textbf{b}_{\gamma})\right\}$ to model the influence of irregular time lags $\delta_{t}$ on the hidden state. Finally, BRITS performs the computation process in both the forward and backward direction so that it can make use of both historical and future information to fill in the missing values.
Loss design: BRITS adopts the reconstruction-based SSL strategy for model training. Formally,
$\mathcal{L}=\mathcal{L_{\textit{e}}}(\bf{x}_{\textit{t}},\hat{\bf{x}}_{\textit{t}})+\mathcal{L_{\textit{e}}}(\bf{x}_{\textit{t}},\hat{\bf{z}}_{\textit{t}})+\mathcal{L_{\textit{e}}}(\bf{x}_{\textit{t}},\hat{\bf{c}}_{\textit{t}}),$ (7)
where $\hat{\textbf{z}}_{t}$ is feature-based estimation, and $\hat{\textbf{c}}_{t}=\beta_{t}\odot\hat{\textbf{z}}_{t}+(1-\beta_{t})\odot\hat{\textbf{x}}_{t}$, $\beta_{t}\in[0,1]$ used as the weight of combining $\hat{\textbf{x}}_{t}$ and $\hat{\textbf{z}}_{t}$. $\mathcal{L}_{\textit{e}}$ is the mean absolute error.
V-B $E^{2}$
E2GAN [34] is a GAN-based end-to-end model for multivariate time series imputation. It leverages an auto-encoder in the generator, where the encoder adds random noise $\eta$ to the input with missing values, compressing it into a low-rank latent vector $z$, the decoder then reconstructs into a complete time series. Both the encoder and decoder utilize GRUI [18] to handle missing values. The discriminator consists of a GRUI layer followed by a fully connected layer to output the probability of the sample being realistic.
Loss design: E2GAN leverages the WGAN [100] training strategy to improve stability and mitigate mode collapse. It incorporates the reconstruction-based SSL to train the generator, enabling the generation of realistic and high-quality data. The discriminator’s loss is defined as:
$\mathcal{L}_{D}=-D(\mathbf{X})+D(\hat{\mathbf{X}})$ (8)
To impute missing data, the generator’s loss incorporates a mean squared error:
$\mathcal{L}_{G}=\lambda\|\mathbf{X}\odot\mathbf{M}-G(\mathbf{X}+\eta)\odot\mathbf{M}\|_{2}^{2}-D(\mathbf{X}^{\prime})$ (9)
To strengthen the generator, E2GAN updates the discriminator every $k$ iteration of generator updates during training.
V-C mTAN
mTAN [101] is a VAE-based time series imputation model. By combining attention and VAE, the sparse and irregularly sampled time series can be projected into a latent space. Unlike conventional attention, mTAN calculates attention scores by time series and a series of reference points. The encoder of mTAN projects irregularly sampled time series data into a latent space of fixed dimension through specially designed attention, and the decoder recovers missing data based on the attention between the latent vector and reference points.
Loss design: mTAN employs the reconstruction-based SSL strategy to train the model. The learning objective of mTANs is shown below, with some minor changes from the VAE,
$\begin{split}\mathcal{L}_{\mathrm{NVAE}}(\theta,\gamma)=\sum_{t=1}^{T}\frac{1}{N}(\mathbb{E}_{q_{\gamma}\left(\mathbf{z}\mid\mathbf{r},\mathbf{x}_{t}\right)}\left[\log p_{\theta}\left(\hat{\mathbf{x}}_{t}\mid\mathbf{z},t\right)\right]\\ -D_{\mathrm{KL}}(q_{\gamma}\left(\mathbf{z}\mid\mathbf{r},\mathbf{x}_{t}\right)\|p(\mathbf{z}))),\end{split}$ (10)
where $\mathbf{z}$ is the defined latent variable assumed to follow a multidimensional normal distribution $p(\mathbf{z})$. $q_{\gamma}$ is the distribution obtained by the encoder, $p_{\theta}\left(\hat{\mathbf{x}}_{t}\mid\mathbf{z},t\right)$ refers to the probability of the decoder obtaining $\hat{\mathbf{x}}_{t}$ based on $\mathbf{z}$ and $t$.
V-D IGNNK
IGNNK[62] focuses on spatial-temporal kriging task, whose structure consists of three graph neural network layers. The most important design of IGNNK is that it can use the message passing mechanism of the graph neural network to aggregate the information of neighbor nodes inductively so as to be applied to other datasets with different graph structures without retraining.
Loss design: IGNNK implements the masked SSL strategy to train the model, yet it simultaneously evaluates observed and missing data. To make the model inductive for unseen graphs, it randomly samples subgraphs in each batch to simulate different graph structures. At each subgraph, IGNNK randomly selects some nodes as the sampled nodes and the others as unsampled nodes. To make the message passing mechanism more generalized for all nodes, IGNNK defines the loss function as follows, which uses the total reconstruction error on both observed and unseen nodes:
$\mathcal{L}=\sum\left\|\hat{\mathbf{X}}-\mathbf{X}\right\|^{2}$ (11)
V-E LATC
LATC [14] is a low-rank autoregressive tensor completion model designed for spatial-temporal traffic data. The model initially reshapes the traffic data matrix $\mathbf{X}\in\mathbb{R}^{N\times T}$ into a third-order tensor $\mathcal{X}\in\mathbb{R}^{N\times I\times J}$ to capture the periodic of traffic data, where $I$ represents the number of time slots within a day, $J$ represents the number of days, and $I\times J=T$. Subsequently, LATC follows to the low-rank assumption, aiming to minimize the tensor rank to effectively capture the global patterns. Meanwhile, LATC introduces a temporal variation term as a new regularization component to capture local temporal consistency. The temporal variation is defined as the cumulative sum of autoregressive errors within the unfolded time series matrix. To approximate the low-rank characteristic and circumvent the challenges associated with determining the rank in factorization models, LATC employs the truncated nuclear norm of the completed tensor.
Loss design: It introduces an autoregressive norm $\left\|\hat{X}\right\|_{W,{\mathcal{H}}}$ to capture local trends within the data. This autoregressive norm is defined as follows, which is a combination of the reconstruction-based SSL loss function and autoregression:
$\left\|\hat{\mathbf{X}}\right\|_{\mathbf{W},{\mathcal{H}}}=\sum_{n=1}^{N}\sum_{t=1}^{T}[\hat{\mathbf{X}}_{n,t}-(\sum_{i}^{d}\mathbf{W}_{n,i}\hat{\mathbf{X}}_{n,t-h_{i}})]^{2},$ (12)
where $d$ is a hyperparameter to determine the number of historical values, $h_{i}$ represents the time interval between the time of the $i$-th historical value and the current time $t$, =$\mathcal{H}=\{h_{1},\cdots,h_{d}\}$ denotes the set of these time intervals, and $\mathbf{W}\in\mathbb{R}^{N\times d}$ is the coefficient matrix that models the contribution of each historical value to the current time step.
Additionally, LATC preserves global trends in the imputed data by introducing a forward tensorization operator $\mathcal{Q}(\cdot)$ that transforms the imputed traffic matrix $\hat{\mathbf{X}}$ into a third-order tensor $\hat{\mathcal{X}}=\mathcal{Q}(\hat{\mathbf{X}})$. To ensure the low-rank characteristic of the tensor and capture the global patterns, LATC minimizes the nuclear norm of the third-order tensor $\|\hat{\mathcal{X}}\|_{r,*}$. Formally,
$\min_{\hat{\mathcal{X}},\hat{\mathbf{X}},W}\left\|\hat{\mathcal{X}}_{r,*}\right\|+\frac{\lambda}{2}\left\|\hat{\mathbf{X}}_{W,\mathcal{H}}\right\|,$ (13)
[…]
where the operation $\mathcal{P}_{\Omega}(\cdot)$ retains the observable values while setting the missing positions to 00, and $r$ is a truncation parameter.
V-F PriSTI
PriSTI [71] is a self-supervised spatiotemporal data imputation model based on diffusion. The model consists of two components: the Conditional Feature Extraction Module (CFEM) and the Noise Estimation Module (NEM). CFEM learns spatiotemporal prior first linearly impute data, then using graph neural networks and attention mechanisms. NEM serves as the noise prediction module within the diffusion framework, employing both temporal and spatial attention to capture spatiotemporal correlations and geographical relationships. CFEM provides NEM with a better global contextual prior knowledge, aiding in its further understanding of spatiotemporal dependencies. During the reverse denoise process, CFEM and NEM progressively recover real data from the noise by multi-step denoise.
Loss design: PriSTI adopts diffusion loss with the masked SSL strategy to evaluate the added noise on missing data. During training, it only computes the error at the locations where missing values are artificially introduced. The loss function is similar to the conventional diffusion model.
V-G GCASTN
GCASTN is a spatial-temporal imputation model that utilizes generative-contrastive learning. It constructs two views to reconstruct the complete data using the same encoder and decoder. In the encoder, GCASTN introduces the $\delta$ to represent the time interval between the current missing position and its closest observable position in the past, then adaptively adjusts the learned spatial-temporal dependencies to make the imputation more reliable.
Loss design: GCASTN adds contrastive learning based on the masked SSL, wherein its loss function comprises two parts. The first is reconstructing the missing values in two views:
$\mathcal{L}_{G}=\sum(\left\|\hat{\mathbf{X}}^{\prime}-{\mathbf{X}}\right\|_{2}^{2}+\left\|\hat{\mathbf{X}}^{\prime\prime}-{\mathbf{X}}\right\|_{2}^{2})$ (14)
The second loss is the contrastive loss that aligns different views:
$\mathcal{L}_{C}=\sum(\left\|\hat{\mathbf{X}}^{\prime}-\hat{\mathbf{X}}^{\prime\prime}\right\|_{2}^{2}),$ (15)
where $\hat{\mathbf{X}}^{\prime}$ and $\hat{\mathbf{X}}^{\prime\prime}$ is the first and second positive sample. The total loss $\mathcal{L}=\mathcal{L}_{G}+\alpha\mathcal{L}_{C}$ by a weight $\alpha$.
V-H AGCRN
AGCRN is a classic graph neural network for traffic prediction. It designs a novel data-driven node embedding, and then implements adaptive GCN through node embedding, thus getting rid of the constraints of prior graph structure. In addition, it reduces the complexity of GCN through matrix decomposition operations while providing an independent parameter space for each node.
Loss design: The original AGCRN uses the mean absolute error to predict future flow. Here, we train it using the masked SSL strategy, making it suitable for imputation tasks.
V-I ASTGNN
ASTGNN is an attention-based spatiotemporal graph neural network designed for traffic forecasting. It utilizes the same sequence-to-sequence structure as the Transformer. To model the local trends within the data, ASTGNN introduces a novel trend-aware attention mechanism that discovers local change trends from the local context. In summary, ASTGNN can effectively capture both global and local spatiotemporal dependencies within the data, thereby enabling accurate forecasting of spatiotemporal data.
Loss design: To enable ASTGNN for the imputation task, we adopt the same masked SSL strategy as AGCRN.
V-J ImputeFormer
ImputeFormer [55] combines the Transformer with low-rank induction to have both prior structure and representation capabilities for spatiotemporal imputation. It balances inductive bias and model expressivity to handle missing data in various settings. The model primarily incorporates attention mechanisms in both the temporal and spatial dimensions, with a special design imposing low-rank constraints on the two attention operations. Specifically, for temporal attention, ImputeFormer applies low-rank constraints by projecting the data into a low-dimensional space. For spatial attention, it replaces the traditional attention score computation with inner products of low-dimensional node embeddings, thereby imposing low-rank constraints.
Loss design: ImputeFormer utilizes masked SSL to train the model. Its loss function integrates MAE loss $\mathcal{L}_{\mathrm{recon}}$ to reconstruct the masked part. It also introduces a Fourier Sparsity Regularization $\mathcal{L}_{\mathrm{FIL}}$ to maintain low-rank structures by imposing spectral constraints on imputations. The loss function is defined as follows:
$\begin{split}\mathcal{L}_{\mathrm{recon}}&=\sum\|\mathbf{M}\odot(\widehat{\mathbf{X}}-\mathbf{X})\|_{1},\\ \mathcal{L}_{\mathrm{FIL}}&=\sum\|\mathrm{Flatten}(\mathrm{FFT}(\bar{\mathbf{X}})\|_{1},\\ \bar{\mathbf{X}}&=\mathbf{M}\odot\mathbf{X}+(\mathbf{1-M})\odot\widehat{\mathbf{X}},\end{split}$ (16)
where FFT means the Fast Fourier Transform.
VI Experiments
VI-A Datasets
We evaluate the representative imputation models mentioned above using public datasets on two kinds of traffic imputation tasks, including highway traffic flow imputation and highway traffic speed imputation. The detailed statistical information of datasets is summarized in Table II.
VI-A1 Highway Traffic Flow
In the highway traffic flow imputation task, we conduct experiments on two commonly used datasets [102], namely PEMS04 and PEMS08, which are collected by the Caltrans Performance Measurement System (PeMS) and the traffic flow is aggregated into 5-minute windows. PEMS04 records 307 sensors traffic flow from January 1st, 2018 to February 28th, 2018, and PEMS08 records 170 sensors traffic flow from July 1st, 2016 to August 31st, 2016.
VI-A2 Highway Traffic Speed
We choose Seattle dataset [103] to evaluate the highway traffic speed imputation task. This dataset is collected in Seattle area, which records the speed of 323 loop detectors from January 1st, 2015 to December 31st, 2015 and the time interval is 1 hour.
TABLE II: Dataset Description.
Data type Datasets # Nodes Time Range Time interval
Traffic Flow PEMS04 307 01/01/2018 - 02/28/2018 5 minutes
PEMS08 170 07/01/2016 - 08/31/2016 5 minutes
Traffic Speed Seattle 323 01/01/2015 - 12/31/2015 1 hour
Figure 3: The illustration of the pipeline, which consists of missing scenario construction, data pre-processing, model construction, and performance evaluation.
VI-B Evaluation Pipeline
We propose a unified evaluation pipeline for traffic data imputation, as illustrated in Figure 3. This pipeline standardizes the whole process of evaluating imputation models by four steps, including missing scenario construction, data pre-processing, model construction, and evaluation.
VI-B1 Missing Scenario Construction
We construct 20 missing scenarios based on real-world conditions, including four distinct missing patterns, each with five missing rates $\alpha$. The construction process of each missing pattern is as follows:
    SRTR. We randomly generate a mask matrix $\mathcal{M}$ with the same dimensions as the original data $\mathcal{X}$. Each element in $\mathcal{M}$ is assigned a value of 0 with a probability of $\alpha$, and a value of 1 with a remaining probability of $1-\alpha$.
    SRTC. In this missing pattern, we initially define a continuous time splice of length $l_{m}$ as a patch and divide the dataset across the time dimension into $\frac{T}{l_{m}}$ non-overlapping patches. We then introduce missing scenarios at the patch level, randomly masking $\frac{T}{l_{m}}*N*\alpha$ patches.
    SCTR. For the spatial continuous missing, we employ a graph clustering algorithm [104] to group all sensors into $N_{c}$ clusters based on distance. And each sensor is assigned a cluster. We construct missing scenarios at the cluster level, randomly masking out $T*N_{c}*\alpha$ clusters.
    SCTC. We initially generate $\frac{T}{l_{m}}*N_{c}$ non-overlapping blocks. Then we introduce missing scenarios at the block level, randomly masking out $\frac{T}{l_{m}}*N_{c}*\alpha$ blocks.
VI-B2 Data Pre-process
The dataset undergoes a three-step pre-processing procedure. Initially, we establish the sliding window size of $T_{w}$, to generate data samples for model training. For the LATC model, all training samples are input simultaneously. Next, acknowledging that many traditional imputation models are specifically designed for multi-variable time series without spatial dimensions, we reshape the data samples to align with each model’s input format. Concurrently, we assign a value of 0 to the missing data locations. Lastly, we apply the normalization method detailed in the original paper to standardize the data.
VI-B3 Model Construction
We summarize 10 recently proposed sequence imputation and prediction models. For constructing the model, we follow the three steps: First, we adapt the code from the official release and only adjust the data input and output to align with traffic data. Secondly, we conduct a large-scale grid search to optimize their hyperparameters for peak performance. Finally, we train each model using a consistent training dataset, guided by a loss function, with an early stopping criterion applied to the validation dataset.
VI-B4 Performance Evaluation
We conduct a detailed evaluation of 20 missing scenarios and 10 baselines across three key aspects. For effectiveness, we use three commonly adopted metrics to assess mean error: Root Mean Square Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE). For efficiency, we compare their training time, inference time, and memory usage. Additionally, we evaluate their imputation performance across both difficult and simple time intervals to simulate varying real-world traffic conditions. Note that our evaluation only focuses on the performance at the locations where data is missing.
VI-C Performance Evaluation Results
VI-C1 Imputation Performance on Different Datasets
We report the imputation performance results of different models on the PEMS04, PEMS08, and Seattle datasets across four missing patterns and five missing rates.
Figure 4: Performance of traffic data imputation for different models with 20 missing scenarios at PEMS04.
Figure 5: Performance of traffic data imputation for different models with 20 missing scenarios at PEMS08.
As shown in Figures 5, 5, and 6, we observe that the performance declines as the missing rate increases. When the missing rate is below 0.5, the performance change of each model remains relatively stable. However, as the missing rate reaches 0.7 and 0.9, the performance of all models significantly deteriorates. This indicates that the noise of missing values at lower missing rates is minimal and does not substantially impact the model. In contrast, when the missing rate exceeds 0.5, the abundance of missing values disrupts the inherent spatial-temporal relationships in the data.
Figure 6: Performance of traffic data imputation for different models with 20 missing scenarios at Seattle.
ImputeFormer, GCASTN, BRITS, and PriSTI are the top-4 best deep-learning models for traffic data imputation, a common characteristic of their specialized design for handling missing data. ImputeFormer incorporates low-rankness into Transformers to capture spatiotemporal structures and mitigate the noise impact of missing value. GCASTN and BRITS use the time decay strategy to introduce time intervals between observed and missing values, while PriSTI applies linear imputation to capture temporal context. Therefore, efficiently incorporating prior knowledge of missing data may be a promising direction in traffic data imputation.
BRITS, E2GAN, mTAN. These three models focus primarily on capturing temporal correlation. In the Seattle dataset, BRITS outperforms E2GAN and mTAN, which can be attributed to its effective time decay strategy that captures small flow changes within local time. BRITS also reports well-performance on time continuous missing patterns in the PEMS04 and PEMS08 datasets, highlighting the suitability of time decay for modeling data trends. mTAN performs well with temporal random missing patterns by capturing global temporal relationships. It is neglected by BRITS and E2GAN, which primarily emphasize local temporal correlations.
IGNNK, PriSTI, GCASTN. These three models account for spatial-temporal correlation. GCASTN excels in high missing rate scenarios (above 50%) due to its contrastive learning approach, emphasizing data quality and model robustness. PriSTI delivers impressive results in low missing rate scenarios since it can deal with the missing value in advance and further consider their spatial-temporal dependence. IGNNK’s underperformance stems from two factors: 1) the sparse adjacency matrice hinders spatial information messages, and 2) the lack of consideration for temporal information.
LATC and ImputeFormer. In contrast to deep learning models, LATC leverages the low-rank assumption of the matrix for model optimization. ImputeFormer combines low-rank with Transformer, further improving the performance. This indicates that taking into account the low-ranking features offers new insights on imputation tasks.
TABLE III: Performance comparison at difficult and simple intervals.
Model Difficult Intervals Simple Intervals
MAE RMSE MAPE MAE RMSE MAPE
LAST 34.02 51.28 13.89 11.42 19.23 25.28
BRITS 25.64 40.95 10.25 10.21 17.22 26.84
E2GAN 33.37 50.46 13.18 11.73 18.42 30.38
IGNNK 26.82 39.48 10.71 8.54 13.34 22.30
mTAN 30.84 48.69 12.05 11.30 18.32 27.36
PriSTI 30.82 46.57 11.89 11.22 17.53 23.64
GCASTN 22.77 34.96 8.79 7.34 11.47 18.30
AGCRN 27.21 39.90 10.51 8.48 13.22 20.36
ASTGNN 26.53 42.62 9.96 8.21 13.22 20.34
LATC 27.89 45.35 11.05 9.80 16.43 27.43
ImputeFormer 24.31 39.35 9.32 7.46 12.52 17.34
VI-C2 Performance During Difficult and Simple Intervals
Most imputation models evaluate average performance in the test set, which fails to capture their strengths and weaknesses across different intervals. For instance, data imputation is more challenging during periods of fluctuating flow changes, while smoother flow transitions are easier to handle. To address this, we sort test samples by variance, with the top 25% with the highest variance classified as difficult intervals and the bottom 25% as simple intervals. We report performance across these intervals using the PEMS04 dataset with a 0.5 missing rate under the SRTR missing pattern, as shown in Table III. The results show that performance decreases significantly in difficult intervals due to frequent fluctuations, whereas simple intervals, characterized by smoother flow, exhibit more predictable temporal relationships.
By comparing models, ImputeFormer and GCASTN demonstrate superior performance in difficult and simple intervals, respectively, highlighting the importance of spatial-temporal modeling. ImputeFormer effectively captures global knowledge and adapts to fluctuations, while GCASTN is well-suited for steady trends by utilizing the time decay strategy. GCASTN also achieves the best and second-best performance in difficult intervals due to its contrastive learning strategy, which enhances robustness. BRITS and E2GAN underperform as they focus solely on temporal dependencies, while spatial-temporal modeling methods such as ASTGNN and AGCRN produce more balanced performance.
Figure 7: Model performance comparison at different missing patterns.
Figure 10: Visualization of model performance across different missing patterns with the missing rate of 0.5 on PEMS04.
VI-C3 Ranks on Model Performance across Different Missing Patterns
The different missing patterns caused the imputation performances of the model to be different. To explore which technology is effective for missing patterns, we count the times of each model that achieved the best, the top-2, and the top-3 performance. If a model has more times, its performance is better. As shown in Figure 7, We observe that only BRITS, GCASTN, PriSTI, ImputeFormer, and LATC achieve the top 3 best performance. A common characteristic of them is introducing prior knowledge, where BRITS, GCASTN, and PriSTI use time delay, and ImputeFormer and LATC consider low-rank characteristics to capture the global correlation. Yet other models only consider the spatio-temporal correlation, failing to efficiently imputation since extra noise of missing data leads to the spatiotemporal correlation being inaccurate. Thus, we guess efficiently extra prior knowledge from data may be a promising direction for more accurate imputation.
TABLE IV: Time and memory cost. / notes that the model does not need to infer stage.
Model Training time (s) Inference time (s) Memory (MB)
BRITS 2511.77 17.46 1612MB
E2GAN 312.42 0.86 1513MB
IGNNK 632.00 1.59 4485MB
mTAN 955.49 4.22 1749MB
PriSTI 68762.25 8553.77 3195MB
GCASTN 138565.63 107.46 7594MB
AGCRN 7013.01 17.49 7579MB
ASTGNN 48539.37 84.74 6961MB
LATC 829.00 / 98MB
ImputeFormer 15320.95 8.19 5234MB
VI-D Efficiency Evaluation Results
We evaluate the training time, inference time, and memory usage of each model on the PEMS04 dataset with the SRTR missing pattern at a 0.5 missing ratio. Notably, the LATC model does not require inference. To ensure consistency, we adhere to the official code for all models, including the number of training epochs and early stopping settings. Detailed results are presented in Table IV.
LATC has the smallest training time and memory usage. E2GAN, IGNNK, and mTAN also require less memory because their components are based on GRU or GNN with small computation costs. GCASTN, ImputeFormer, and ASTGNN utilize transformer architecture, so they demand more memory due to the quadratic space complexity of attention mechanisms. PriSTI exhibits significantly longer inference times. This is attributed to the diffusion model, which involves iterative denoising during inference and lacks parallelization support. Additionally, the need for multiple evaluations to estimate uncertainty further increases its inference time.
VI-E Case study
To more intuitively observe the difference between the imputation results of each model and the ground-truth, we conduct a case study for each missing pattern with a 0.5 missing rate at two different type datasets PEMS04 and Seattle.
For the spatial random missing pattern, we randomly select a node to visualize the impute results and ground truth. As shown in Figure 10 (a) and (b), we observe that the overall imputation curve of each model is aligned with the ground truth, demonstrating their ability to capture the temporal dependence. Among them, BRITS, GCASTN, ImputeFormer, and LATC are more closely to real data. A common characteristic of these models is their incorporation of prior knowledge about missing data, which enhances imputation accuracy. BRITS and GCASTN leverage time-delay strategies to model temporal dependencies, while LATC and ImputeFormer adopt low-rank approaches to better capture the underlying data structure.
For the spatial continuous missing patterns, we select a cluster of missing nodes, visualize their connection relationships and observe the MAE metric of each node. The results as shown in Figure 10 (c) and (d). At SCTR missing pattern, ASTGNN, GCASTN, BRITS, and PriSTI have smaller errors, indicating the importance of modeling temporal dependencies, especially when spatial information is unavailable. We also notice that IGNNK performs poorly under the SCTC missing pattern due to its only focus on local spatial correlations and ignoring temporal dependencies, making it ineffective for this type of missing pattern.
VI-F Summary of New Insights
Through a large of experimental analysis, we find some interesting insights.
VI-F1 Effective Components
(1) Incorporating prior knowledge to initialize missing values is important. To initialize missing values, BRITS and GCASTN employ the time delay mechanism, and PriSTI adopts the linear imputation mechanism. And they achieve more accurate imputation performance. Compared to them, IGNNK and mTAN use 0-initiation, exhibiting unsatisfactory performance. (2) Tensor completion, which captures global patterns like periodicity in traffic data, is useful. We can see that TC based ImputeFormer and LATC outperform most deep learning models in many scenarios.
VI-F2 Model Choosing in Various Scenarios
(1) Different missing patterns. GCASTN is the best model for temporal continuous missing pattern, and BRITS is the best model for spatial continuous missing pattern. (2) Memory-constrained scenarios. TC is the best choice since LATC has the smallest memory cost. Among deep learning methods, the diffusion model is the most appropriate choice for balancing performance and memory. (3) Scenarios with real-time requirements for inference speed. TC has the fastest inference speed. For deep learning methods, E2GAN is the fastest, and BRITS is the second.
VI-F3 Relations between Prediction and Imputation
We find that spatial-temporal graph prediction models ASTGNN and AGCRN can achieve comparable performance in traffic imputation tasks, particularly in low missing rate scenarios. We think the reasons for this phenomenon is both of these two kinds of models focus on capturing the dynamics of traffic data along the spatial and temporal dimensions. Meanwhile, we can find that with the increase of missing rate, prediction models fail. This may because prediction model do not distinguish missing values and observed values, i.e., only treat them as different numbers. With the increase of missing rate, missing values may misguide models. Thus, as long as we can design a suitable mechanism to indicate missing values or use a proper missing initializing mechanism, it is expected to unify these two tasks.
VII Conclusion
In this paper, we propose a practice-oriented categorization of imputation models. We classify the real traffic missing scenarios into four missing patterns and five missing rates. The imputation models are further organized based on their spatial-temporal modeling techniques and loss function designs. Additionally, we propose a unified pipeline to evaluate these models across three standardized traffic datasets and comprehensive missing patterns and analyze their performance in terms of efficiency, effectiveness, and different time intervals. Through extensive experimental analysis, we draw some valuable insights among effective components, model choosing in various scenarios, and the relation between imputation and prediction, which provide valuable guidance for future research.